OpenStudy (anonymous):

20. A cylindrical vase has a diameter of 6 inches. At the bottom of the vase, there are 6 marbles, each of diameter 1 inch. The vase is filled with water up to a height of 12 inches. What is the volume of water in the vase?

OpenStudy (campbell_st):

find the volume of the cylinder, but first get the radius. \[V = \pi r^2 h \] find the volume of 3 spherical marbles... but first find the radius of the marbles. \[V = 3 \times \frac{4}{3} \pi r^3\] the volume of water will be Volume of the cylinder x the volume of 3 marbles.

OpenStudy (campbell_st):

oops Volume of Water = Volume of Cylinder - Volume of the marbles... sorry about the typo

OpenStudy (anonymous):

thank you i got it now :)